OpenStudy (anonymous):

Given the data points (1, 4), (4, 13), and (6, 19), which of the following equations describes the best-fit line?

OpenStudy (anonymous):

it would help to know the following equations

OpenStudy (anonymous):

A. = 1 + 3x B. = 3 + 1x C. = 2 + 3x D. = -1 + 3x

OpenStudy (anonymous):

possibly C but dont take my word on it

OpenStudy (imstuck):

Those points are linear, so finding the equation simply requires that you find the slope and where the line goes through the y axis. The line goes through the y axis at y = 1. So 1 is your y intercept. Then find the slope.

OpenStudy (imstuck):

The slope goes up 3 and over 1, so slope is 3. In y = mx + b, the m is the slope and the b is the y intercept, both of which we know. So the equation is y = 3x + 1, or 1 + 3x in your answer choices, which is A.

OpenStudy (anonymous):

its was A thank you @IMStuck and thank for trying @malibu597